The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by checkley71(8403)

4B-8=ONE OF THE EQUAL LEGS.
THUS THE PERIMETER IS:
2(4B-8)+B=29
8B-16+B=29
9B=29+16
9B=45
B=45/9
B=5 ANSWER FOR THE LENGTH OF THE BASE.
4*5-8=12 FOR EACH OF THE EQUAL LEGS.
PROOF
2*12+5=29
24+5=29
29=29